# A Third Paradigm for LLM Evaluation: Dialogue Game-Based Evaluation using clembench

### v1

@@ A Third Paradigm for LLM Evaluation:
David Schlangen1,2   Sherzod Hakimov1   Jonathan Jordan1   Philipp Sadler1 1 Computational Linguistics, University of Potsdam, Germany 2 DFKI (German Research Center for AI), Berlin
@@ 1 Introduction
[…]
@@
When a new large language model is released, it is customary now to provide two types of evaluation results (see, e.g., Team et al. (2025)): First, results are given from user rankings, as collected on the LM Arena (formerly Chatbot Arena, Chiang et al. (2024)). This is meant to represent general user perception of model quality. Second, results on “standard benchmarks” are provided, for example such as are collected by the LLM-harness Gao et al. (2023). This, in turn, is meant to measure specific capabilities in more depth and with more control than is brought out by the user preferences. While these two types of evaluation complement each other (controlled and static, less controlled and interactive), it has recently been argued that a third combination (controlled and interactive) is needed to acquire a well-rounded impression of the capabilities of a model Momentè et al. (2025).
As discussed below (Section 2.3), there have been several attempts at realising controlled, interactive evaluation through making LLMs self-play conversational games. Arguably, widespread adoption of this paradigm has been held back by the fact that many of these attempts appear to have been built for a single evaluation campaign only, with less attention being paid to re-usability of code. The package highlighted here, clembench Chalamalasetti et al. (2023), was one of the first instantiations of this idea, and has driven a leaderboard that has been continuously updated since early 2023. Thanks to a grant by the University of Potsdam, we can now offer clembench2, which has been optimized for ease of use in various use cases, including local evaluation of LLMs or LLM fine-tunes. In this paper, we first briefly describe the “dialogue game-based evaluation” paradigm in general, explaining how it differs from the other paradigms (as illustrated in the feature matrix in Table 1). We then focus specifically on clembench2, with the use case of evaluating new model in mind. For this, we show how the existing benchmark can be run, to yield results that can directly be compared to the long history of measurements recorded on the official leaderboard.222 https://clembench.github.io We also show how easily the benchmark can be extended with games highlighting special abilities.
@@ 2.1 The General Idea
To address this question, we can see that even this simple interaction challenges certain capabilities: understanding of game rules and how they constrain the available actions; understanding of a target concept sufficiently to be able to describe it; ability to construct utterances under negative constraints (avoiding certain lexical items); decoding a clue; ability to integrate repeated clues. As has been argued in the literature discussed below, these are capabilities that static datasets such as MMLU-Pro Wang et al. (2024), Big-bench Hard Suzgun et al. (2023), and IFEval Zhou et al. (2023) do not, or at least not as systematically and holistically, test.333 See also Schlangen (2023, 2019) for an extensive discussion of this question.
@@ 2.2 Some Unique Properties
What makes dialogue game-based evaluation (DGBe) special? For Table 1, we have compiled into a feature matrix some of the properties as claimed in the work cited below. As can be seen, it is the combination of properties that makes DGBe special. Through the selection of games, instantiations, and game instances as non-changing, pre-constructed elements, a high degree of control over the task is conferred, while the fact that by self-play, only models and no human players are involved, provides a certain degree of replicability. Unlike in static datasets, many of the games allow for or even require multi-turn interactions. The separation of game and game instance makes it possible to explore different ways of presenting game material; the clembench, for example, provides, for some games, coordinated text-only and text/image variants Hakimov et al. (2025). The fact that pre-defined experiments are being evaluated also brings DGBe closer to reference-based evaluation (RBe), and makes the cost of running the benchmark on a model more predictable and fixed; whereas in preference-based evaluation (PBe), this in principle is unlimited and the results get better / more stable, the longer the model is provided and compared. The last three features mentioned in Table 1 all relate to the relative ease with which new game instances and even new games can be created (more on that below), compared to static datasets, which need to be authored or collected and curated.
@@ 2.3 Existing Frameworks
The clem framework demonstrated in this paper has two components: A) clemcore, which is the pip-installable backbone abstracting access to LLM into a common interface and providing the infrastructure for setting up games and running experiments on them (e.g., logging, scoring, aggregating results). B) clembench, which is a collection of implemented games, and, for each version number of the benchmark, a different pre-compiled set of instances to run.555clemcore: https://github.com/clembench/clemcore; clembench: https://github.com/clembench/clembench. Both are licensed under the permissive MIT license. This division of labour is illustrated in Figure 3.
The framework has been used to power the public ‘clembench leader board’ since 2023.666https://clembench.github.io/leaderboard.html The long period for which this has already been run has made it possible to track interesting trends; for example, the plot in Figure 5 (Appendix) clearly shows the moment at which open-weight models started to catch up, as well as the generally fast improvements over time. (See Beyer et al. (2024) for a more detailed discussion.) We also host a transcript browser which makes it particularly easy to browse the by now quite large database of self-play transcripts of various types of proprietary and open-weight models. [^7] Transcript browser: https://clembench.github.io/transcript-browser.html. The transcripts are also all stored at https://github.com/clembench/clembench-runs/ (more than 3GB by now).
@@ 4.2 Creating New Games
4.3 Adding New Languages
As explained, the way a game is instantiated so that LLMs can be made to play it is via textual descriptions on the one hand, and the game logic implemented in code on the other. The latter part is language agnostic. The clembench as referred to above comes with prompts written in English, and, for games where this is relevant (e.g., wordle, taboo), with instances in English. While this is something we have only begun to explore ourselves (see initial results reported by Beyer et al. (2024)), the framework is ready for defining experiments with stimuli in other languages. The flexibility of the machinery for running experiments (see Section 3) makes it possible to easily define language-specific as well as cross-lingual benchmark runs. We are currently setting up a mechanism for leveraging the community to collect game instantiations in a large set of languages.
@@ 5 Conclusions
This paper accompanies a demonstration of the clembench2 evaluation package. We hope this paper has illustrated that running game-based evaluations does not need to be complicated, and in fact, using clembench, in most cases is a matter of running a single command (and providing a single file, describing the model thamt is to be evaluated). The reasons why models should be evaluated in this way and what exactly is being measured could only be sketched here, and for a full argumentation, we refer the reader to recent works such as Momentè et al. (2025) as well as the papers mentioned in Section 2.3 above. We hope that the work reported here can contribute to further adoption of this complementary evaluation paradigm, in particular in light of the increased use of LLMs as backbones for interactive and agentic tasks.
Appendix A Existing Frameworks and Their Maintenance Status
Table 2 lists all related frameworks. The column “Maintained” is determined whether there have been changes on the respective Git repository in the last three months.
@@ Appendix A Existing Frameworks and Their Maintenance Status
Appendix B Trends Observed on the clembench leader board
See Figure 5.

### v2

@@ A Third Paradigm for LLM Evaluation:
David Schlangen1,2   Sherzod Hakimov1   Chalamalasetti Kranti1 Jonathan Jordan1   Philipp Sadler1 1 Computational Linguistics, University of Potsdam, Germany 2 DFKI (German Research Center for AI), Berlin
@@ 1 Introduction
leakage danger $\uparrow$ $\downarrow\downarrow$ $\downarrow$
extension cost (instances) $\uparrow$ open $\downarrow\downarrow$
@@
When a new large language model is released, it is customary now to provide two types of evaluation results (see, e.g., Gemma Team et al. (2025)): First, results are given from user rankings, as collected on the LM Arena (formerly Chatbot Arena, Chiang et al. (2024)). This is meant to represent general user perception of model quality. Second, results on “standard benchmarks” are provided, for example such as are collected by the LLM-harness Gao et al. (2023). This, in turn, is meant to measure specific capabilities in more depth and with more control than is brought out by the user preferences. While these two types of evaluation complement each other (controlled and static, less controlled and interactive), it has recently been argued that a third combination (controlled and interactive) is needed to acquire a well-rounded impression of the capabilities of a model Momentè et al. (2025).
As discussed below (Section 2.3), there have been several attempts at realising controlled, interactive evaluation through making LLMs self-play conversational games. Arguably, widespread adoption of this paradigm has been held back by the fact that many of these attempts appear to have been built for a single evaluation campaign only, with less attention being paid to re-usability of code. The package highlighted here, clembench Chalamalasetti et al. (2023), was one of the first instantiations of this idea, and has driven a leaderboard that has been continuously updated since early 2023. Thanks to a grant by the University of Potsdam, we can now offer clembench2, which has been optimized for ease of use in various use cases, including local evaluation of LLMs or LLM fine-tunes. In this paper, we first briefly describe the “dialogue game-based evaluation” paradigm in general, explaining how it differs from the other paradigms (as illustrated in the feature matrix in Table 1). We then focus specifically on clembench2, with the use case of evaluating new model in mind. For this, we show how the existing benchmark can be run, to yield results that can directly be compared to the long history of measurements recorded on the official leaderboard. [^2] https://clembench.github.io/leaderboard.html We also show how easily the benchmark can be extended with games highlighting special abilities. [^3] The scientific basis behind the approach is discussed, evaluated, and validated at length in Chalamalasetti et al. (2023); Hakimov et al. (2025a). The use of the framework as an environment for reinforcement learning of multi-turn problems has been explored by Horst et al. (2025). The present paper aims at making the community aware of it as a resource for easy evaluation of new models or model checkpoints during development.
@@ 2.1 The General Idea
To address this question, we can see that even this simple interaction challenges certain capabilities: understanding of game rules and how they constrain the available actions; understanding of a target concept sufficiently to be able to describe it; ability to construct utterances under negative constraints (avoiding certain lexical items); decoding a clue; ability to integrate repeated clues. As has been argued in the literature discussed below, these are capabilities that static datasets such as MMLU-Pro Wang et al. (2024), Big-bench Hard Suzgun et al. (2023), and IFEval Zhou et al. (2023) do not, or at least not as systematically and holistically, test. [^4] See also Schlangen (2023, 2019) for an extensive discussion of this question.
@@ 2.2 Some Unique Properties
What makes dialogue game-based evaluation (DGBe) special? For Table 1, we have compiled into a feature matrix some of the properties as claimed in the work cited below. As can be seen, it is the combination of properties that makes DGBe special. Through the selection of games, instantiations, and game instances as non-changing, pre-constructed elements, a high degree of control over the task is conferred, while the fact that by self-play, only models and no human players are involved, provides a certain degree of replicability. Unlike in static datasets, many of the games allow for or even require multi-turn interactions. The separation of game and game instance makes it possible to explore different ways of presenting game material; the clembench, for example, provides, for some games, coordinated text-only and text/image variants Hakimov et al. (2025a). The fact that pre-defined experiments are being evaluated also brings DGBe closer to reference-based evaluation (RBe), and makes the cost of running the benchmark on a model more predictable and fixed; whereas in preference-based evaluation (PBe), this in principle is unlimited and the results get better / more stable, the longer the model is provided and compared. The last three features mentioned in Table 1 all relate to the relative ease with which new game instances and even new games can be created (more on that below), compared to static datasets, which need to be authored or collected and curated.
@@ 2.3 Existing Frameworks
The clem framework demonstrated in this paper has two components: A) clemcore, which is the pip-installable backbone abstracting access to LLM into a common interface and providing the infrastructure for setting up games and running experiments on them (e.g., logging, scoring, aggregating results). B) clembench, which is a collection of implemented games, and, for each version number of the benchmark, a different pre-compiled set of instances to run.666clemcore: https://github.com/clembench/clemcore; clembench: https://github.com/clembench/clembench. Both are licensed under the permissive MIT license. This division of labour is illustrated in Figure 3.
The framework has been used to power the public ‘clembench leader board’ since 2023.777https://clembench.github.io/leaderboard.html The long period for which this has already been run has made it possible to track interesting trends; for example, the plot in Figure 5 (Appendix) clearly shows the moment at which open-weight models started to catch up, as well as the generally fast improvements over time. (See Beyer et al. (2024) for a more detailed discussion.) We also host a transcript browser which makes it particularly easy to browse the by now quite large database of self-play transcripts of various types of proprietary and open-weight models. [^8] Transcript browser: https://clembench.github.io/transcript-browser.html. The transcripts are also all stored at https://github.com/clembench/clembench-runs/ (more than 3GB by now).
Currently, 15 games with text-only input and 6 multimodal games (image and text inputs) are made publicly available with more being developed in the pipeline.
@@ 4.2 Creating New Games
4.3 Multilingual Evaluation with clembench: Adding New Languages
As explained, the way a game is instantiated so that LLMs can be made to play it is via textual descriptions on the one hand, and the game logic implemented in code on the other. The latter part is language agnostic. The clembench as referred to above comes with prompts written in English, and, for games where this is relevant (e.g., wordle, taboo), with instances in English. While this is something we have only begun to explore ourselves (see initial results reported by Beyer et al. (2024)), the framework is ready for defining experiments with stimuli in other languages. Hakimov et al. (2025b) evaluated the negotiation abilities and multilingual aspect in LLMs by developing dialogue games in three languages: English, German, Italian. The flexibility of the machinery for running experiments (see Section 3) makes it possible to easily define language-specific as well as cross-lingual benchmark runs. We are currently setting up a mechanism for leveraging the community to collect game instantiations in a large set of languages.
5 “Vibe-checking” the Agentic Abilities of a Model
A new addition in clembench2 is the integration of the dialogue experiment framework slurk Götze et al. (2022). This integration makes it possible to interact locally (on one’s own machine) with a model as a partner in the implemented dialogue games and thus to get a feel for how it behaves in constrained agentic conditions. [^14] A common practice that has been systematically argued for e.g. by Dunlap et al. (2025). The slurk framework has also proven to be able to handle larger crowdsourcing experiments (e.g., Ilinykh et al. (2019)), so that more thorough quantitative studies are also possible. [^15] A preview of the public interface is hosted at https://clemp.ling.uni-potsdam.de; note however that we cannot guarantee that backend models as game play partners are available at any time, so this should be seen only as a demo of the interface. The game play itself will be demonstrated at the conference.
6 Evaluating the Evaluator
6.1 Human Game-play Performance
To assess the validity of games and to compare with LLM performance, in Beyer et al. (2024) we performed a user study on the initial set of games. We asked the participants to play 10-15 episodes per game (leaving out wordle-clue and wordle-critic, as these are only variants of the main wordle game). All episodes were played to the end where the human participants obliged the formatting-related game rules (reaching 100% in Played score). The resulting Quality scores were as follows: wordle: 72, taboo: 80.5; drawing: 95.2; reference: 100; leading to an average of 86.93. The best performing model o3-mini on the clembench leaderboard (v2.0) still lags behind the average human performance, it has an average Quality score of 82.23 on the same games.
6.2 Reproducibility of Benchmark Results
All games used for benchmarking are publicly available as well as the their specific instance versions. We run the vanilla (without any agentic or other tools) version of benchmarked LLMs by keeping the same parameters (temperature, library/API version, etc.) across them. Thus, as long as the model providers do not make changes to their models (applies to commercial models), the benchmark results are reproducible. (To the extent that LLM calls are deterministic. However, slight differences in responses are unlikely to affect final scores, which are an average over the perfomance on a larger set of instances.) All existing benchmark runs are made publicly available. [^16] https://github.com/clembench/clembench-runs
6.2.1 The Value of Game-Based Evaluation
Momentè et al. (2025) have compared the use of games for interactive evaluation (as implemented in clembench), the use of targetted cognitive test sets, and the use of more generalist benchmarking datasets. A central finding was that games provide a more fine-grained means to discriminate between models; the authors recommend an approach that “triangulates” performance through a combination of interactive and static benchmarking.
@@ 7 Conclusions
This paper accompanies a demonstration of the clembench2 evaluation package. We hope this paper has illustrated that running game-based evaluations does not need to be complicated, and in fact, using clembench, in most cases is a matter of running a single command (and providing a single file, describing the model that is to be evaluated). The reasons why models should be evaluated in this way and what exactly is being measured could only be sketched here, and for a full argumentation, we refer the reader to recent works such as Momentè et al. (2025) as well as the papers mentioned in Section 2.3 above. We hope that the work reported here can contribute to further adoption of this complementary evaluation paradigm, in particular in light of the increased use of LLMs as backbones for interactive and agentic tasks.
Appendix A Appendix
Table 2 lists all related frameworks. The column “Maintained” is determined whether there have been changes on the respective Git repository in the last three months. Figure 5 shows trends observed on the clembench leader board. Figure 6 shows an example transcript.